Solve the following pair of linear equations by the substitution method
 3x – y = 3
9x – 3y = 9 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given, 3x – y = 3
and
9x – 3y = 9
are the two equations.
From 1st equation,
we get,
x = (3+y)/3
Now, substitute the value of x in the given second equation to get,
9(3+y)/3 – 3y = 9
⇒9 +3y -3y = 9
⇒ 9 = 9
Therefore, y has infinite values
and since,
x = (3+y) /3,
so x also has infinite values.

Thanks